OpenStudy (anonymous):

If 18 meters of high-resistance wire with a diameter of 1.2 mm has a resistance of 10 ohms, what is the resistance of 27 m of the same type of wire with a diameter of 1.5 mm?

OpenStudy (irishboy123):

use \[R = \frac{\rho L}{A}\]

OpenStudy (anonymous):

ok?

OpenStudy (irishboy123):

if rho (resistivity) is the same, then \[\left( \frac {RA}{L} \right)_1 = \left( \frac {RA}{L} \right)_2\]

OpenStudy (radar):

. The resistance of the wire will vary directly with length and inversely with area. In this case I would guess the resistance of the wire will increase but not by much